An analysis is carried out on a sample pf unknown gas. The density of the gas is 2.86 g/L at STP. The molecular formula is

A.HCL

B. Cl2

C.NO2

D.SO2

can you plz provide an explanation for this plz?
1
Expert's answer
2017-02-18T04:20:12-0500
As it is known, the volume of one mole of any gas at STP is 22.41 liters. Also it is known that the mass of one mole of any substance is equal to its molecular weight. Knowing the density of our gas sample we can find the mass of one mole of this gas, and it will be its molecular weight.

Mass = volume * density.
Calculation:
Mass = 22.41 L * 2.86 g/L = 64.1 g.
So the molecular weight of our unknown gas is 64.1. Now we can check molecular weights (Mr) of all proposed substances and see which one fits.
a) HCl: Mr = 1.0 + 35.5 = 36.5. No.
b) Cl2: Mr = 35.5*2 = 71.0. No.
c) NO2: Mr = 14.0 +16.0*2 = 46.0. No.
d) SO2: Mr = 32.1 + 16.0*2 = 64.1. Yes!

Answer:
The unknown gas is SO2.